The equation -2x+5y=3 goes through which point and has what slope?
Answer by checkley71(8403)

-2X+5Y=3 OR 5Y=2X+3 OR Y=2/5X+3/5 THUS THE SLOPE OF THIS LINE IS 2/5 THE Y INTERCEPT IS 3/5 AND THEA POINT THROUGH WHICH THE LINE PASSES IS (O,3/5) &
(-3/2,0)